The length of arectangular flower bed is 3 feet less than twice its width. the area of the bed is 54 square feet. find the dimensions of the flower bed?
Answer by edjones(8007)

L=2w-3
Lw=54
.
w(2w-3)=54 substitute 2w-3 for L
2w^2-3w-54=0
2*-54=-108. -12 and 9 are the factors of -108 whose sum is -3.
2w^2-12w+9w-54=0
2w(w-6)+9(w-6)=0
(2w+9)(w-6)=0
.
2w=-9
w=-9/2 answer can't be negative.
.
w=6'
L=9'
.
Ed